i can't figure out the formula if x = -1 ; and y = -1/2
can you please help me.
also another one is x = -2; and y = -9
Thank you

Answer by stanbon(75887)

i can't figure out the formula if x = -1 ; and y = -1/2
can you please help me.
also another one is x = -2; and y = -9
-------------------
If you are looking for the equation of the line that
passes thru (-1, -1/2) and (-2,-9), do the following:
--------------
slope = (-9 - (-1/2))/(-2--1) =(-17/2)/(-1) = 17/2
-------------
Formula is y = mx+b and you now know y,x,and m; solve for b as follows:
-9 = (17/2)(-2) + b
-9 = (-17) + b
b = 8
-----------------
EQUATION:
y = (17/2)x + 8
======================
Cheers,
Stan H.